OpenStudy (gigiheart63):

help equation down below what is the radius Write your answer in simplified, rationalized form.

OpenStudy (gigiheart63):

\[\left( x+\frac{ 15 }{ 2 } \right)^{2}+y^2=36\]

satellite73 (satellite73):

what are you supposed to do with it?

OpenStudy (gigiheart63):

what is the radius

satellite73 (satellite73):

the square root of the number on the right of the equal sign

satellite73 (satellite73):

general form of circle with center \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (gigiheart63):

@JoshMorris

OpenStudy (jtm-pa):

Ms brains